You have been managing a $5 million portfolio that has a beta of 0.85 and a required rate of return of 7.975%. The current risk-free rate is 5%. Assume that you receive another $500,000. If you invest the money in a stock with a beta of 0.55, what will be the required return on your $5.5 million portfolio? Do not round intermediate calculations. Round your answer to two decimal places.

______ %
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

Curently:

7.975=5+0.85*(market rate-5)

(7.975-5)=0.85*(market rate-5)

market rate=(7.975-5)/0.85+5

=8.5%

Hence required rate for $500,000=5+0.55*(8.5-5)

=6.925%

Required rate of portfolio=Respective rate*Respective weight

=(5,000,000/5,500,000*7.975)+(500,000/5,500,000*6.925)

=7.88%(Approx).